# CNNCat: Categorizing high-energy photons in a Compton/Pair Telescope with Convolutional Neural Networks

### v1

@@ 1 Introduction
Gamma-ray astronomy in the energy range of sub-MeV to several 10s of MeV remains little explored. At these energies, the Compton effect dominates the total cross-section of photons with matter, changing to dominant pair production in the upper part of this range. The intransparency of the atmosphere requires Compton and pair telescopes to be operated from space. Currently, INTEGRAL 2003A&A…411L…1W provides varying sensitivity from 15 keV to 10 MeV, while Fermi-LAT Atwood_2009 covers the range from 30 MeV to 300 GeV. Higher energy photons are observed by ground-based observatories like H.E.S.S. osti_20875888 , Veritas veritas and MAGIC magic or, in the future, CTA cta . More than 20-year old all-sky maps in 0.7 – 30 MeV are available from the Compton telescope COMPTEL comptel on the Compton Gamma-Ray Observatory (CGRO), yet at the limited sensitivity of a pioneering instrument. Comparatively poor sensitivity in the energy range from about 0.1 MeV to 100 MeV leaves a window open for discovery, often referred to as the MeV Gap.
@@ 1.2 Misclassification and Data Handling in Space
Bandwidth in Low Earth Orbit is a limited resource. It has to be shared with housekeeping data, satellite telemetry, and secondary instruments on the respective satellite bus. One important goal of this study therefore consists in reliably identifying and discarding events that will not be usable in a later stage of event reconstruction. Event classification is crucial in reaching this goal. A small and computationally lean model that divides events into potentially reconstructable ”good” events and non-reconstructable or background events can help to reduce the amount of data that has to be transmitted. It can also improve event-type-specific preprocessing onboard. Application of an incorrect reconstruction algorithm to an otherwise valid event leads to a false reconstruction of source photons into the background. The purpose of this study is to investigate whether it is possible to classify events with high efficiency and speed into Compton events, reconstructable Pair events, and non-reconstructable events (3-class classification), while restricting the evaluation of the pursuing improvement to Compton data only.
@@ 1.3 Deep Feedforward Convolutional Neural Networks
where $\mathbf{x}$ is an input vector and $\mathbf{\Theta}$ are the learned model parameters Goodfellow-et-al-2016 . In the case of a classification task, $\mathbf{y}$ is identified with a category. Feedforward means that data is processed inside the model only in one direction (towards the output nodes). There are no feedback connections that loop data back into previous layers of the model.
Convolutional Neural Nets (CNNs) CNNs emerged as the go-to processing approach when dealing with images and gridded data with local dependencies. The main idea is to restrict the search for relevant information to small regions of the input by applying a moving filter with learned kernel values on the pixel values. The number of free (trainable) parameters of a pure CNN is independent of the input size and lower than in any conventional Multi Layered Perceptron (MLP) approach since kernel sizes in general do not exceed $7\times 7=49$ entries. This is much smaller than the actual number of pixels in an image, which would form the input for an MLP. The number of calculations necessary in a CNN depends on the size of the input. To scale this down, a common strategy is to sample down the intermediate input by either maximum or average pooling: applying a moving window to the input, either the maximum or the average value of the pixels inside the window is propagated.
@@ 2.1 Architecture
A preliminary investigation using the approach presented by Aurisano_2016 led to the conclusion that the use of their Inception modules would lead to a rather slow and computationally expensive model. The SqueezeNet architecture DBLP:journals/corr/IandolaMAHDK16 , however, enables the design of small, yet powerful models, which have been demonstrated to run on small-scale hardware like FPGA-accelerated System-On-Chips (SoCs) zynqNet and was chosen as design pattern of the convolutional modules. The presented model makes heavy use of SqueezeNet’s Fire modules shown in the top right of Fig. 3 and is sketched in Fig. 3 left.
@@ 2.2 Dataset Generation and Selection Criteria
It was also deemed important to enforce clean samples from which the model is tasked to learn the characteristics of good Compton and Pair events. The very first interaction of a photon in each training set sample, independent of the type of event, has to occur in sensitive detector material. Pair events have to be created inside the tracker. Since vetoed events in space-borne detectors are usually discarded immediately, they are not considered in the datasets of this study. For image augmentation, we applied Gaussian noise with the estimated energy resolution of the proposed detector to the input images to further solidify the model against overfitting. [^1] Since the D1 detector is highly inhomogenious due to its layered design, the common technique of rotating the image cannot be applied.
@@ 3.1 Computational Demands and Architectural Optimizations
Applying this to the precursor model allowed us to reduce the number of trainable parameters to 64k and the number of FLOPS per event to $0.9$ GFLOPS. ESA identified the necessity to explore systems for onboard processing and funded the GPU4S (GPUs for Space) project that surveyed use-cases and possible embedded systems GPU4S . Follow-up studies also investigated potential improvements of commercial off-the-shelf embedded systems (e.g. GPUFollowUp ). Ortiz et al. also studied the Jetson AGX system which delivers around 1.3 TFLOPS per Watt at FP16 or 2.6 TOPS at INT8, equivalent to a rate of 1.4 kHz or 2.9  kHz per Watt, respectively. Note that either a smaller detector or a coarser binning of x-y strips in the input data would increase the amount of rate per Watt significantly as the computational complexity for CNNs scales mostly with $j\times k$ of the input data. According to NVIDIA’s claims, space-graded systems are to be released. Therefore, a model similar to the one presented here is suited to run on soon to be available embedded systems.
@@ 3.2 Performance Measures
It gives the fraction of correct predictions relative to all predictions. Accuracy is a workable metric on class-wise balanced datasets but may give deceptive results when the data is skewed towards one class. For example, consider a case where the signal-to-background ratio is expected to be very small. A (useless) ’classifier’ that would tag all events as background would reach an accuracy of almost 100% while failing the given task completely. On unbalanced datasets, performance must be measured with more robust metrics. Two of those, from the context of binary classification and information retrieval, are precision and recall. Precision is defined as
@@
Recall, also called true positive rate, is a measure of how well a given class is recognized by a classifier, i.e., the efficiency with which these events are properly classified. In general, it is impossible to optimize both precision and recall to arbitrarily high values unless the samples in the dataset are perfectly separable. Usually, precision and recall reach an equilibrium state where improvements in one metric imply performance loss in the other.
Both metrics can be extended from binary to multi-class scenarios by applying a one-against-all scheme. The class to be evaluated is considered the positive class, while the other classes are combined into the negative class. False positives then have to be defined as the combined misclassifications into the investigated class from all other classes. Table 2 shows the confusion matrix for the 3-class task to be performed.
Table 2: Layout of the confusion matrix for the 3-class task in this study. True positives align on the main diagonal. Each column corresponds to all samples predicted into a given class, each row provides the classification result for a given (true) class. C is for Compton event, P is for pair event, NR is for non-reconstructable event. Every off-diagonal matrix element can be interpreted as either a false negative or a false positive of a given class.
[…]
Compton tp C fn C, fp P fn C, fp NR
Pair fn P, fp C tp P fn P, fp NR
non-reconstructable fn NR, fp C fn NR., fp P tp NR
@@ 3.3 Training and Validation
The training set consisted of 400,000 events evenly divided over classes, incident energy and incident angle binning. 300k stratified events were used for training, 100k similar selected events were held back for validation for the purpose of monitoring the training progress. We used the Adam optimizer kingma2017adam for parameter adaptation. Early stopping was used to prevent overfitting due to excessive training. The progress during training is shown in Figure 5. The relevant training and hyper parameters are summarized in Table 3
@@ 3.4 Classification Performance
To assess the total performance of our classifier, we calculated the Receiver-Operator-Characteristic Area-under-the Curve (ROC-AUC) by determining the True Positive Rate and False Positive Rate for decreasing decision thresholds for the positive class. In multi-class scenarios this is done in a so called ‘one vs. all’ approach. There is only one positive class and the other classes are merged into a single negative class. This score can be used to assess the performance of a classifier. The resulting curves for this test dataset are shown in Fig. 9. All three classes are well separated from random guessing (indicated by the dashed line). There is no task-independent interpretation of the score, but one can employ a rule of thumb found in rocAuc . Performance could be considered ‘poor’ for a score below 0.7, ‘acceptable’ for 0.7 to 0.8, ‘excellent’ for 0.8 to 0.9, and ‘outstanding’ for scores larger than 0.9. The scores for each class are around a value of 0.8, which indicates a good performance on the presented dataset. The best performance is achieved by the pair classification, which achieves an area of 0.9. According to the cited performance standards, our model reaches the second highest performance category.
@@ 3.5 Crab Nebula Simulation
In this part, we investigate the impact of prior classification on the reconstruction of Compton event data. To do so, we applied the standard MEGAlib reconstruction chain on a simulated Crab Nebula measurement. The Crab spectrum was modelled according to ref. crab_spectrum for energies up to 6 MeV, and extended to 20 MeV by a power-law with index 2.24, as shown in Fig. 10. The Crab simulation was performed in eight separate energy ranges. For each bin, a simulation with 500k pre-trigger events was performed, resulting in a total of 2.85 M events with a distribution shown in Table 4. The source was set up as a far-field point source with an incident angle $\theta=30^{\circ}$ MEGAlib was tasked with reconstructing Compton and Pair events and eliminating events from minimum ionizing particle tracks (MIPs), using the standard reconstruction chain.
@@
MEGAlib performs its default event classification at the reconstruction stage in a cascade from MIP to pair to Compton recoil track search, and finally to Compton sequence reconstruction. We considered an event tagged as Compton when the reconstruction succeeded in Compton reconstruction. The event was considered tagged as pair, when the pair reconstruction accepted the event. An event was considered tagged as non-reconstructable when it was either tagged as MIP or all reconstruction algorithms rejected the event.
@@
Figure 12: Precision and Recall for Pair and non-reconstructable events as function of energy for the Crab-source dataset.
The performance on D1 and D2 pairs is depicted in Figure 12. At energies above the pair threshold and below 5 MeV, both the model and the default classification show very little discrimination power. Nonetheless, the model’s Recall on D2-first pairs shows an improvement over the default case. The D1-pair Recall shows large improvements over the default classification, starting to recognize D1-pairs reliably at energies larger than 5 MeV and a very good class-coverage of 87 % at the highest simulated energies. The Precision of the model classifications on D1 and D2-pairs shows a slightly worse purity of the pair datasets after classification than in the default case, which can be explained by the relatively large amount of incorrectly tagged D2-pairs which have been tagged as D1-pair and vice-versa (see Figs. 6, 6 and 6).

### v2

@@ 1 Introduction
Gamma-ray astronomy in the energy range of sub-MeV to several 10s of MeV remains little explored. At these energies, the Compton effect dominates the total cross-section of photons with matter, changing to dominant pair production in the upper part of this range. The intransparency of the atmosphere requires Compton and pair telescopes to be operated from space. Currently, INTEGRAL 2003A&A…411L…1W provides varying sensitivity from 15 keV to 10 MeV, while Fermi-LAT Atwood_2009 covers the range from 30 MeV to 300 GeV. Higher energy photons are observed by ground-based observatories like H.E.S.S. hess , Veritas veritas and MAGIC magic or, in the future, CTA cta . More than 20-year old all-sky maps in 0.7 – 30 MeV are available from the Compton telescope COMPTEL comptel on the Compton Gamma-Ray Observatory (CGRO), yet at the limited sensitivity of a pioneering instrument. Comparatively poor sensitivity in the energy range from about 0.1 MeV to 100 MeV leaves a window open for discovery, often referred to as the MeV Gap.
@@ 1.2 Misclassification and Data Handling in Space
Bandwidth in Low Earth Orbit is a limited resource. It has to be shared with housekeeping data, satellite telemetry, and secondary instruments on the respective satellite bus. One important goal of this study therefore consists in reliably identifying and discarding events that will not be usable in a later stage of event reconstruction. Event classification is crucial in reaching this goal. A small and computationally lean model that divides events into potentially reconstructable ”good” events and non-reconstructable or background events can help to reduce the amount of data that has to be transmitted. It can also improve event-type-specific preprocessing onboard. Application of an incorrect reconstruction algorithm to an otherwise valid event leads to a false reconstruction of source photons into the background. The purpose of this study is to investigate whether it is possible to classify events with high efficiency and speed into Compton events, reconstructable Pair events, and non-reconstructable events (3-class classification), while restricting the evaluation of the pursuing improvement to energies up to 20 MeV.
@@ 1.3 Deep Feedforward Convolutional Neural Networks
where $\mathbf{x}$ is an input vector and $\mathbf{\Theta}$ are the learned model parameters Goodfellow-et-al-2016 . In the case of a classification task, $\mathbf{y}$ is identified with a category. Feedforward means that data is processed inside the model only in one direction towards the output nodes.
Convolutional Neural Nets (CNNs) CNNs emerged as the go-to processing approach when dealing with images and gridded data with local dependencies. The main idea is to restrict the search for relevant information to small regions of the input by applying a moving filter with learned kernel values on the pixel values. The number of free (trainable) parameters of a pure CNN is independent of the input size and lower than in any conventional Multi Layered Perceptron (MLP) approach since kernel sizes in general do not exceed $7\times 7=49$ entries. This is much smaller than the actual number of pixels in an image, which would form the input for an MLP. The number of calculations necessary in a CNN depends on the size of the input. To scale down the processing overhead, a common strategy is to sample down the intermediate input by applying either maximum or average pooling. Here a moving window is applied to the input, propagating either the maximum or the average value of the pixels inside the window.
@@ 2.1 Architecture
The SqueezeNet architecture DBLP:journals/corr/IandolaMAHDK16 , enables the design of small, yet powerful models, which have been demonstrated to run on small-scale hardware like FPGA-accelerated System-On-Chips (SoCs) zynqNet and was chosen as design pattern of the convolutional modules. The presented model makes heavy use of SqueezeNet’s Fire modules shown in the top right of Fig. 3 and is sketched in Fig. 3 left.
@@ 2.2 Dataset Generation and Selection Criteria
It was also deemed important to enforce clean samples from which the model is tasked to learn the characteristics of good Compton and Pair events. The very first interaction of a photon in each training set sample, independent of the type of event, has to occur in sensitive detector material. Pair events have to be created inside the tracker. Since vetoed events in space-borne detectors are usually discarded immediately, they are not considered in the datasets of this study. For image augmentation, we applied Gaussian noise with the estimated energy resolution of the proposed detector to the input images to further solidify the model against overfitting. [^1] Since the D1 detector is highly inhomogeneous due to its layered design, the common technique of rotating the image cannot be applied.
@@ 3.1 Computational Demands and Architectural Optimizations
Applying this to the precursor model allowed us to reduce the number of trainable parameters to 64k and the number of FLOPS per event to $0.9$ GFLOPS. ESA identified the necessity to explore systems for onboard processing and funded the GPU4S (GPUs for Space) project that surveyed use-cases and possible embedded systems GPU4S . Follow-up studies also investigated potential improvements of commercial off-the-shelf embedded systems (e.g. GPUFollowUp ). Ortiz et al. also studied the Jetson AGX system which delivers around 1.3 TFLOPS per Watt at FP16 or 2.6 TOPS at INT8, equivalent to a rate of 1.4 kHz or 2.9  kHz per Watt, respectively. Note that either a smaller detector or a coarser binning of x-y strips in the input data would increase the amount of rate per Watt significantly as the computational complexity for CNNs scales mostly with $j\times k$ of the input data. A more recent study of Rodriguez-Ferrandez et al in cooperation with ESA was able to show that this kind of devices can sustain the radiation environment in orbit XavierProton . They obtained acceptable error rates, making them a candidate system for embedded systems especially designed for space applications.
@@ 3.2 Performance Measures
It gives the fraction of correct predictions relative to all predictions. Accuracy is a workable metric on class-wise balanced datasets but may give deceptive results when the data is skewed towards one class. On unbalanced datasets, performance must be measured with more robust metrics. Two of those, from the context of binary classification and information retrieval, are precision and recall. Precision is defined as
@@
Recall, also called true positive rate, is a measure of how well a given class is recognized by a classifier, i.e., the efficiency with which these events are properly classified. In general, it is impossible to optimize both precision and recall to arbitrarily high values unless the samples in the dataset are perfectly separable. Both metrics can be extended from binary to multi-class scenarios by applying a one-against-all scheme. The class to be evaluated is considered the positive class, while the other classes are combined into the negative class. False positives then have to be defined as the combined misclassifications into the investigated class from all other classes.
@@ 3.3 Training and Validation
The training set consisted of 400,000 events evenly divided over classes, incident energy and incident angle binning. 300k stratified events were used for training, 100k similar selected events were held back for validation for the purpose of monitoring the training progress. We used the Adam optimizer kingma2017adam for parameter adaptation. Early stopping was used to prevent overfitting due to excessive training. During trials we observed that the improvement stalled after only a couple of epochs. To mitigate this, early stopping was used to trigger a restart of the optimizer with a lower learn-rate on the best-performing model (denoted as Training Reset) in order to force the optimizer to a lower minimum. The progress during training is shown in Figure 5. The relevant training and hyper parameters are summarized in Table 2
@@ 3.4 Classification Performance
To assess the total performance of our classifier, we calculated the Receiver-Operator-Characteristic Area-under-the Curve (ROC-AUC) by determining the True Positive Rate and False Positive Rate for decreasing decision thresholds for the positive class. In multi-class scenarios this is done in a so called ‘one vs. all’ approach. The resulting curves for the test dataset are shown in Fig. 9. All three classes are well separated from random guessing (indicated by the dashed line). There is no task-independent interpretation of the AUC-score, but one can employ a rule of thumb found in rocAuc . The scores for each class are around a value of 0.8, which indicates a good performance on the presented dataset. The best performance is achieved by the pair classification, which achieves an area of 0.9. Following the scheme of rocAuc , our model reaches the second highest performance category.
@@ 3.5 Crab Nebula Simulation
In this part, we investigate the impact of prior classification on the reconstruction of event data up to 20 MeV, applying the standard MEGAlib reconstruction chain on a simulated Crab Nebula measurement. The Crab spectrum was modelled according to ref. crab_spectrum for energies up to 6 MeV, and extended to 20 MeV by a power-law with index 2.24, as shown in Fig. 10. The Crab simulation was performed in eight separate energy ranges. For each bin, a simulation with 500k pre-trigger events was performed, resulting in a total of 2.85 M events with a distribution shown in Table 3. The source was set up as a far-field point source with an incident angle $\theta=30^{\circ}$ MEGAlib was tasked with reconstructing Compton and Pair events and eliminating events from minimum ionizing particle tracks (MIPs). It applies a cascade of reconstruction algorithms with increasing computational complexity to each event. If at any stage the algorithm is successful, the chain stops and the event is classified accordingly. The MIP-search tests the hit-pattern for linearity by applying a linear $\chi^{2}$ fit. The next stage is the Pair-reconstruction. Here a pattern-recognition algorithm is searching for a creation vertex. If successful, the electron and positron tracks are reconstructed by applying a Pearson correlation test. The Compton reconstruction is separated into two stages. First, the recoil electron track is extracted and reconstructed based on a Spearman-Rank test. Then the Compton-sequence is reconstructed by minimizing the $\chi^{2}$ of the difference between the kinematic scatter angle obtained via Compton’s formula and the geometric scatter angle derived from the hits’ locations. Details about the implementation in MEGAlib can be found inZogDiss .
@@
We use the result of the reconstruction as prediction of the event type of the baseline classification. We considered an event tagged as Compton when the reconstruction succeeded in Compton reconstruction. The event was considered tagged as pair, when the pair reconstruction accepted the event. An event was considered tagged as non-reconstructable when it was either tagged as MIP or all reconstruction algorithms rejected the event.
@@
Figure 12: Precision and Recall for Pair events as function of energy for the Crab-source dataset.
Figure 13: Precision and Recall for non-reconstructable events as function of energy for the Crab-source dataset.
The performance on D1 and D2 pairs is depicted in Figures 12 and 13 . At energies above the pair threshold and below 5 MeV, both the model and the default classification show very little discrimination power. Nonetheless, the model’s Recall on D2-first pairs shows an improvement over the default case. The D1-pair Recall shows large improvements over the default classification, starting to recognize D1-pairs reliably at energies larger than 5 MeV and a very good class-coverage of 87 % at the highest simulated energies. The Precision of the model classifications on D1 and D2-pairs shows a slightly worse purity of the pair datasets after classification than in the default case, which can be explained by the relatively large amount of incorrectly tagged D2-pairs which have been tagged as D1-pair and vice-versa (see Figs. 6, 6 and 6).
@@ 4 Summary and Outlook
Acknowledgement
Jan Peter Lommler acknowledges support by DLR grant 50OO2218.